OpenStudy (anonymous):

How many grams of glucose is in 50.0 mL of a 20% glucose solution mixed with 50.0 mL of water?

OpenStudy (anonymous):

20% = 1/5 so 20% of 50.0 mL = 50.0/5 = 10.0 mL this is then diluted by another 50 mL or a factor of two so 10.0mL / 2 = 5.0mL

OpenStudy (anonymous):

so 5 grams? (because 1g=1mL)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thank you for the explanation:) (my lesson was quite vague)